# Prompt diphoton production at ATLAS 13 TeV in $k_{t}$-factorization: a comparative analysis of UPDF models

### v1

@@ I Introduction
Diphoton production has been extensively studied in various experiments across different kinematic regions and center-of-mass energies [13, 14, 15, 16, 17, 18]. The challenge of achieving satisfactory agreement between theoretical predictions and experimental data, particularly through the inclusion of higher-order subprocesses, has been investigated in numerous studies within the collinear factorization framework [1, 19, 3, 2]. This challenge becomes even more pronounced at higher energies, where next-to-leading order (NLO) Monte Carlo generators such as Diphox [10] and Resbos [11] fail to fully reproduce the observed data.
@@
The $k_{t}$-factorization approach has shown great promise in addressing these challenges [6, 4, 5], offering a more flexible description of initial-state parton dynamics by explicitly accounting for transverse momentum degrees of freedom. Unlike fixed-order collinear calculations, the $k_{t}$-factorization formalism inherently includes a portion of higher-order real emission effects through the intrinsic transverse momentum of the incoming partons [20]. This makes it particularly effective for describing observables sensitive to soft and collinear radiation, such as the azimuthal decorrelation and transverse momentum of the diphoton system.
@@ II.2 The MRW UPDFs method
Martin et al. [34] further demonstrated that using LO splitting functions yields results comparable to those obtained with NLO functions, simplifying the computation. We therefore adopt this approximation, and the NLO-MRW formalism can be expressed as:
@@ III Numerical Methods and Experimental Cuts
    where $E_{T}^{\mathrm{iso}}$ is the sum of transverse energies of the particles around the photon at a distance less than $\Delta R$. The function $\chi(\Delta R)$ is given by:
@@
    Additionally, similar to the simulation with fixed-order NNLO using Nnlojet in Ref. [18], we set $R_{0}=0.1$, $n=2$, and $\epsilon_{\gamma}=0.1$.
@@ IV Results and Discussion
Figure 1: Differential cross sections as functions of the partonic kinematic variables: (left) momentum fractions $x_{1}$ and $x_{2}$, (middle) transverse momenta $k_{t1}$ and $k_{t2}$, and (right) factorization scale $\mu_{\mathrm{fac}}$. The results correspond to the subprocess $q+\overline{q}\to\gamma+\gamma$ at $\sqrt{s}=13~\mathrm{TeV}$.
Figure 2: Comparison of different gluon UPDF models at $\mu=70~\mathrm{GeV}$ (top) and $\mu=500~\mathrm{GeV}$ (bottom). The ratios $PB/MKMRW$, PB/NLOMRWPB/NLO-MRW, and MKMRW/NLOMRWMKMRW/NLO-MRW are shown for each case.
Figure 3: Comparison of different up-quark UPDF models at $\mu=70~\mathrm{GeV}$ (top) and $\mu=500~\mathrm{GeV}$ (bottom). The ratios $PB/MKMRW$, PB/NLOMRWPB/NLO-MRW, and MKMRW/NLOMRWMKMRW/NLO-MRW are shown for each case.
@@
Before discussing and comparing the results obtained with different UPDF models and their impact on the predicted differential cross sections, we first examine the regions of the unintegrated phase space that are most relevant to this study. For this purpose, the Les Houches Event (LHE) files generated by the KaTie parton-level event generator are used to extract the distributions of key partonic variables, and the analysis of the LHE files is performed using the pylhe library [56]. The differential cross sections are studied as functions of the longitudinal momentum fractions $x_{1}$ and $x_{2}$ (the momentum fractions of the incoming partons moving along the $\hat{z}$ and $\hat{z}$ directions, respectively), the transverse momenta $k_{t1}$ and $k_{t2}$ of the initial-state partons, and the factorization scale $\mu_{\mathrm{fac}}$ of the subprocess. These distributions are shown in the figure 1.
In this analysis, we consider two subprocesses, $q+\overline{q}\to\gamma+\gamma$ and $q+g\to\gamma+\gamma+q$, as representative cases. Other subprocesses and UPDF models exhibit qualitatively similar behaviour and are therefore omitted for brevity. As evident from figure 1, a wide range of incoming parton momentum fractions contribute to the calculation, spanning approximately $0.0003\leq x\leq 0.3$. The transverse momenta and factorization scales of the incoming partons also cover large ranges, with the dominant contributions arising from regions where $k_{t}\approx 0.2$$5~\mathrm{GeV}$ and $\mu_{\mathrm{fac}}\approx 70$$150~\mathrm{GeV}$. It should be noted that only a limited number of events from the LHE files are used in this analysis; hence, as visible in figure 1, statistical uncertainties are significant in certain regions, particularly those sensitive to large transverse momenta.

### v2

# Prompt diphoton production compared to measurements at 13 TeV in $k_{t}$-factorization: a comparative analysis of UPDF models

@@ I Introduction
Diphoton production has been extensively studied in various experiments across different kinematic regions and center-of-mass energies [13, 14, 15, 16, 17, 18]. The challenge of achieving satisfactory agreement between theoretical predictions and experimental data, particularly through the inclusion of higher-order subprocesses, has been investigated in numerous studies within the collinear factorization framework [1, 19, 3, 2]. This challenge becomes even more pronounced at higher energies, where next-to-leading order (NLO) calculations such as Diphox [10] and Resbos [11] fail to fully reproduce the observed data.
@@
The $k_{t}$-factorization approach has shown great promise in addressing these challenges [6, 4, 5], offering a more flexible description of initial-state parton dynamics by explicitly accounting for transverse momentum degrees of freedom. Unlike fixed-order collinear calculations, the $k_{t}$-factorization formalism inherently includes a portion of higher-order real emission effects through the transverse momentum of the incoming partons [20]. This makes it particularly effective for describing observables sensitive to soft and collinear radiation, such as the azimuthal decorrelation and transverse momentum of the diphoton system.
@@ II.2 The MRW UPDFs method
Martin et al. [34] further demonstrate that the NLO corrections to the splitting functions are numerically small compared to the LO contributions, and that UPDFs obtained with NLO collinear PDFs but only LO splitting functions reproduce, to very good accuracy, those obtained with the full NLO splitting functions over a wide range of $x$, $k_{t}$ and scales. On this basis they advocate a simplified NLO prescription in which one keeps the NLO PDFs, kinematics and Sudakov form factors, but uses LO splitting functions in the final evolution step. We therefore adopt this simplified NLO-MRW prescription, and the NLO-MRW formalism can be expressed as:
@@ III Numerical Methods and Experimental Cuts
    where $E_{T}^{\mathrm{iso}}$ is the sum of transverse energies of the particles around the photon at a distance less than $\Delta R$. The function $\chi(\Delta R)$ is defined as:
@@
    Following the setup of the Sherpa 2.2 and NNLOJET predictions compared to the same ATLAS measurement [18], we choose
    $R_{0}=0.1,\qquad n=2,\qquad\epsilon_{\gamma}=0.1.$ (25)
    It should be noted that in the ATLAS fiducial phase space a standard cone of radius $\Delta R=0.2$ is used with an isolation requirement $E_{T,\gamma}^{\mathrm{iso},0.2}<0.09\,p_{T,\gamma}$ [18]. For such tight isolation, dedicated studies have shown that smooth-cone and standard-cone prescriptions yield very similar predictions for LHC diphoton observables [56], within theoretical uncertainties, while smooth-cone isolation retains the advantage of suppressing the fragmentation component.
@@ IV Results and Discussion
Figure 1: Comparison of different gluon UPDF models at $\mu=70~\mathrm{GeV}$ (top) and $\mu=500~\mathrm{GeV}$ (bottom). The colour scale shows the ratio of UPDFs as a function of the longitudinal momentum fraction $x$ of the gluon (horizontal axis) and its transverse momentum $k_{T}$ in GeV (vertical axis), both on logarithmic scales. From left to right the panels display the ratios $PB/MKMRW$, $PB/NLO\text{-}MRW$, and $MKMRW/NLO\text{-}MRW$
Figure 2: Comparison of different up-quark UPDF models at $\mu=70~\mathrm{GeV}$ (top) and $\mu=500~\mathrm{GeV}$ (bottom). The colour scale shows the ratio of UPDFs as a function of the longitudinal momentum fraction $x$ of the up quark (horizontal axis) and its transverse momentum $k_{T}$ in GeV (vertical axis), both on logarithmic scales. From left to right the panels display the ratios $PB/MKMRW$, $PB/NLO\text{-}MRW$, and $MKMRW/NLO\text{-}MRW$.
@@
Before discussing and comparing the results obtained with different UPDF models and their impact on the predicted differential cross sections, we briefly comment on the regions of the unintegrated phase space that are most relevant to this study. For this purpose, the Les Houches Event (LHE) files generated by the KaTie parton-level event generator are analysed using the pylhe library [57]. We study the differential cross section as a function of the longitudinal momentum fractions $x_{1}$ and $x_{2}$ of the incoming partons (moving along the $\hat{z}$ and $\hat{z}$ directions, respectively), the transverse momenta $k_{t1}$ and $k_{t2}$ of the initial-state partons, and the factorization scale $\mu_{\mathrm{fac}}$ of the subprocess.
From this analysis we find that, under the ATLAS fiducial cuts, the dominant contribution to the diphoton cross section arises from the ranges $3\times 10^{-4}\lesssim x_{1,2}\lesssim 3\times 10^{-1}$, $0.2~\mathrm{GeV}\lesssim k_{t1,2}\lesssim 5~\mathrm{GeV}$ and $\mu_{\mathrm{fac}}\sim 70$$150~\mathrm{GeV}$. We have verified that the representative subprocesses $q+\overline{q}\to\gamma+\gamma$ and $q+g\to\gamma+\gamma+q$ populate very similar regions of this unintegrated phase space; other channels behave analogously. These ranges are motivated the choices of scales used in the UPDF comparisons shown in Figs. 1 and 2.
@@
Before presenting our predictions for the measured observables, we note that the ATLAS data points shown in this section are taken from the HEPData tables corresponding to Ref. [18]. For each bin, the experimental uncertainty shown in our figures is the total experimental uncertainty, obtained by combining in quadrature all individual components listed in HEPData (statistical and systematic, with upward and downward variations treated separately).